Project

General

Profile

Actions

20160223 » History » Revision 12

« Previous | Revision 12/13 (diff) | Next »
Blackmore, Victoria, 23 February 2016 10:31


Ops/Comm Scheduling meeting

Location : CR11 R3 @ 15:30 pm on Tuesday 23rd

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda

  • Actions from last meeting
    • Pierrick/Mark: Ensure that the pressure and material in the absorber volumes are recorded
    • Mark: Make more accurate measurements of tracker patch panel helium leak
    • Pierrick: Progress on monitoring O2 content of tracker with sensors. Monitoring He levels at TOFs?
    • John G: Resin test progress? (Re: tracker helium leak)
    • Durga/John N: What is the effect on the beamline of removing GVA1?
    • Pierrick: Feedback to alarm handler re: corrupted data was showing false positives
    • Pierrick: Shared services (e.g. water, air) to be included in alarm handler
    • Durga: Convergence on data quality flags needed
    • Durga: Progress on testing of reconstruction quality flags
  • DC report - activities in the Hall this week
  • MOM report - Data-taking this week
    • Lessons learnt during startup
    • Data plots from recent shifts
    • Summary of runs (MOM)
  • Decay Solenoid PS return and calibration
    • Kaeser compressor status
  • Focus coil and absorber installation
  • Controls and monitoring
    • Beamstop summary shows beamstop as closed when it's open
    • Radiation monitor EPICS display freezes for ~2min
    • Proton absorber switch readout doesn't update
    • Q5 and Q6 temperature readback still swapped
  • Scheduling of required hall work and operations
  • List of things to address by February run
  • AOB
    • Surveying of TOF1/TOF2/KL/EMR
    • Tracker waveguide improvements
  • Chair of meetings for next 5 weeks
    • Steve is teaching at this time for the next 4 weeks. DC will chair the meeting.
  • DONM : 1/3/2016

Updated by Blackmore, Victoria over 8 years ago · 12 revisions